STYLE FILES FOR ASCE-LIKE DOCUMENTS

Matthew R. Kuhn^{1}, Member, ASCE

ABSTRACT

This document was produced with the L^{A}TEX typesetting program using the document class “ascelike.cls” and the example document “ascexmpl.tex”. The reference section on page 6 was produced with the BibTEX style “ascelike.bst” and the database “ascelike.bib”. The objective of these files is manuscripts that roughly comply with the guidelines of the American Society of Civil Engineers. The document class produces either double-spaced manuscripts for journal submissions or cameraready manuscripts for conference proceedings. This document serves as a brief guide to ascelike.cls, as well as a test of the output that is produced by the input file ascexmpl.tex. The package is freely available under the LaTeX Project Public License, version 1.1

Keywords: L^{A}TEX, ASCE, document class.

INTRODUCTION

The document class “ascelike.cls” produces manuscripts that roughly comply with the guidelines of the American Society of Civil Engineers (ASCE). The ascelike.cls document class, the bibliographic style ascelike.bst, and example files are available on the ctan web-site (Kuhn 2011). Although it is not produced by ASCE, its agents, or employees, ascelike.cls is now referenced on the ASCE web-site.

The program ascelike.cls is distributed under the terms of the LaTeX Project Public License Distributed, available from the CTAN archives; either version 1.1 of the License, or any later version. If you modify ascelike.cls, you should rename it so that “altered” copies are not later proliferated.

The document class “ascelike.cls” requires the following supplementary

packages: ifthen.sty, setspace.sty, and endfloat.sty Without these files,

ascelike.cls won’t work. These files are typically included in L^{A}TEX distribution, such as the TexLive and MikTex distributions. All of these files are also freely available from the Comprehensive TEX Archive Network (CTAN) archive, through

1Dept. of Civil Engrg., Donald P. Shiley School of Engrg., Univ. of Portland, 5000 N.

Willamette Blvd., Portland, OR 97203. E-mail: kuhn@up.edu.

http://www.ctan.org or http://www.tug.org, although they may need to be

unbundled from a *.dtx form.

In addition to ascelike.cls, the file ascelike.bst can be used with the bibliographic tool BibTEX to produce ASCE-like reference citations and entries (with the weird use of quotation marks around titles, etc.) (Kuhn 2011). An example bibliographic data base is given in ascexmpl.bib.

In addition to these essential files, we have found the following packages very useful (and we will use this opportunity to illustrate a L^{A}TEX itemized list):

• graphicx.sty and its companion files for incorporating encapsulated postscript (figure) files into the document

• times.sty for typesetting with Times fonts.

• subfigure.sty for arranging and numbering sub-figures

• amsmath.sty and its companion files for the AMS extensions to mathe-

      matical formatting (amsfonts.sty, amssymb.sty, and amsbsy.sty).

All of these packages are also freely available from the CTAN archive, but they are included in most L^{A}TEX distributions.

INPUT AND OPTIONS

You should prepare your *.tex input file as a regular L^{A}TEX file using the standard article.cls constructs, but, of course, substitute ascelike for article in the opening \documentclass command. You will likely need to specify a number of options as described below. In addition, ascelike provides two new commands: KeyWords and NameTag, both of which are described further below.

Document class ascelike.cls provides the options given below (and we also use this opportunity to illustrate an enumerated list). The Proceedings|Journal options are the most important; the other options are largely incidental.

1. Options Journal|Proceedings specify the overall format of the output

manuscript.

Journal produces double-spaced manuscripts for ASCE journals. It places tables and figures at the end of the manuscript, and produces lists of tables and figures. It numbers the appendices with Roman numerals and produces proper headings for sections, subsections, subsubsections, appendices, and abstract. It produces the proper page margins and numbers the pages.

Proceedings produces camera-ready single-spaced manuscripts for most ASCE conference proceedings. It produces the proper page margins as given on the old shiny, camera-ready paper (with the light blue lines) supplied by ASCE. It places figures and tables within the text. It produces proper headings for sections, subsections, subsubsections, appendices, and the abstract. Pages are numbered, and the bottom left corner can be “tagged” with the author’s name (this can be done by inserting the command \NameTag{<your name>} within the preamble of your document.

2. Options BackFigs|InsideFigs can be used to override the default placement of tables and figures in the Journal and Proceedings formats.

3. Options SingleSpace|DoubleSpace can be used to override the default text spacing in the Journal and Proceedings formats.

4. Options 10pt|11pt|12pt can be used to override the default text size (12pt).

5. The option NoLists suppresses inclusion of lists of tables and figures that would normally be included in the Journal format.

6. The option NoPageNumbers suppresses the printing of page numbers.

7. The option SectionNumbers produces an automatic numbering of sections. Without the SectionNumbers option, sections will not be numbered, as this seems to be the usual formatting in ASCE journals (note that the appendices will, however, be automatically “numbered” with Roman numerals). With the SectionNumbers option, sections and subsections are numbered with Arabic numerals (e.g. 2, 2.1, etc.), but subsubsection headings will not be numbered. To change this default depth of numbering when the option SectionNumbers is invoked, insert the following commands in the preamble of your document:

   \setcounter{secnumdepth}{1} Number sections only
   \setcounter{secnumdepth}{3} Number sections, subsections,

and subsubsections

Even with the SectionNumbers option, you can use the “starred” form, \section*{ }, to create a section heading without numbers. This might be desirable for an Acknowledgements section at the end of a paper. Note, however, that the starred form will not suppress the numbering of subsections or subsubsections.

SECTIONS, SUBSECTIONS, EQUATIONS, ETC.

This section is included to explain and to test the formating of sections, subsections, subsubsections, equations, tables, and figures. Section heads are automatically made uppercase, which is great unless your section heading contains mathematics, $<math stuff>$. If a heading does contain mathematics, you will need to modify ascelike.cls, in particular the line containing the \uppercase command. To force mathematics symbols to become bold within a section heading, try using the \boldmath command before the in-line math: for example,

\boldmath$a_{i}=\sqrt{\beta}$.

An Example Subsection with math, a_{i} = \sqrt{β}

No automatic capitalization occurs with subsection headings; you will need to capitalize the first letter of each word, as in “An Example Subsection.”

An example subsubsection

No automatic capitalization occurs with subsubsections; you will need to capitalize only the first letter of subsubsection headings. And now we include an FIG. 1. An example figure (just a box). This particular figure has a caption with more information than the figure itself, a very poor practice indeed.

TABLE 1. An example table

Assembly Attribute  Values
(1)  (2)
Number of particles 4008
Particle sizes  Multiple
Particle size range  0.45D_{50}^{∗} to 1.40D_{50}
Initial void ratio, e_{init} 0.179
Assembly size  54D_{50} × 54D_{50} × 54D_{50}

∗ D_{50} represents the median particle diameter

example of a displayed equation (Eq. 1)

E = mc^{2} ,  (1)

a figure (Fig. 1), and a table (Table 1).

A new command \KeyWords{<your key words>} can be used to produce a labeled list of key words. It can be placed anywhere in the document and produces an unindented paragraph of keywords at that location.

CITATIONS AND BIBLIOGRAPHIC ENTRIES

When used together, ascelike.cls and ascelike.bst produce APA / Chicago Manual of Style citations in name-date format. The code in ascelike.bst is a modification of the chicago.sty and chicago.bst packages. The following citation options are available:

• \cite{key} produces citations with full author list and year (Ireland 1954).

• \citeNP{key} produces citations with full author list and year, but without enclosing parentheses: e.g. Ireland 1954.

• \citeA{key} produces citations with only the full author list: e.g. (Ireland)

• \citeN{key} produces citations with the full author list and year, but which can be used as nouns in a sentence; no parentheses appear around the author names, but only around the year: e.g. Ireland (1954) states that …

• \citeyear{key} produces the year information only, within parentheses, as in (1954).

• \citeyearNP{key} produces the year information only, as in 1954.

The bibliographic data base ascexmpl.bib gives examples of bibliographic entries for different document types. These entries are from the canonical set in the ASCE web document “Instructions For Preparation Of Electronic Manuscripts” and from the ASCE web-site. The References section of this document has been automatically created with the ascelike.bst style for the following entries:

• a book (Goossens et al. 1994),

• an anonymous book (Moody 1988),

• an anonymous report using @MANUAL (Federal 1991),

• a journal article (Stahl et al. 2004; Pennoni 1992),

• a journal article in press (Dasgupta 2008),

• an article in an edited book using @INCOLLECTION (Zadeh 1981),

• a building code using @MANUAL (International 1988),

• a discussion of an @ARTICLE (Vesilind 1992),

• a masters thesis using @MASTERSTHESIS (Sotiropulos 1991),

• a doctoral thesis using @PHDTHESIS (Chang 1987),

• a paper in a foreign journal (Ireland 1954),

• a paper in a proceedings using @INPROCEEDINGS (Eshenaur et al. 1991; Garrett 2003),

• a standard using @INCOLLECTION (ASTM 1991),

• a translated book (Melan 1913),

• a two-part paper (Frater and Packer 1992a; Frater and Packer 1992b),

• a university report using @TECHREPORT (Duan et al. 1990),

• an untitled item in the Federal Register using @MANUAL (Federal 1988),

• works in a foreign language (Duvant and Lions 1972; Reiffenstuhl 1982),

• software using @MANUAL (Lotus 1985),

• two works by the same author in the same year (Gaspar and Koenders 2001a; Gaspar and Koenders 2001b), and

• two works by three authors in the same year that only share the first two authors (Huang et al. 2009a; Huang et al. 2009b).

ASCE has added two types of bibliographic entries: web-pages and CD-ROMs. A web-page can be formated using the @MISC entry category, as with the item (Burka 1993) produced with the following *.bib entry:

 @MISC{Burka:1993a,
   author = {Burka, L. P.},
   title = {A hypertext history of multi-user dimensions},
   journal = {MUD history},

      year = {1993},
      month = {Dec. 5, 1994},
      url = {http://www.ccs.neu.edu}
    }

Notice the use of the “month” field to give the date that material was downloaded and the use of a new “url” field. The “url” and month” fields can also be used

with other entry types (i.e., @BOOK, @INPROCEEDINGS, @MANUAL, @MASTERSTHESIS,

@PHDTHESIS, and @TECHREPORT): for example, in the entry type @PHDTHESIS for (Wichtmann 2005).

A CD-ROM can be referenced when using the @BOOK, @INBOOK, @INCOLLECTION, or @INPROCEEDINGS categories, as in the entry (Liggett and Caughey 1998). The field “howpublished” is used to designate the medium in the .bib file:

    howpublished = {CD-ROM},

MISCELLANY

Many ASCE conference proceedings are now published on CD ROM media. I have noticed that instructions on paper formats issued by conference organizers often differ from the standard ASCE instructions. Fortunately most of the differences can be easily accommodated, such as changes in the margins and placement of the authors’ addresses. As for margins, these can, of course, be altered by using \setlength{<length>} commands within the preamble to a document without making any changes to ascelike.cls. (See the L^{A}TEX book (Lamport 1994), its companion (Goossens et al. 1994), or online web documentation.)

Authors’ addresses can be placed below the title (instead of in a footnote) by not using the \thanks command.

REFERENCES ASTM (1991). “Standard practice for the use of the international system of units

(SI) (the modernized metric system).” E 380-91a, ASTM, Philadelphia, Pa. Burka, L. P. (1993). “A hypertext history of multi-user dimensions.” MUD his-

tory, <http://www.ccs.neu.edu> (Dec. 5, 1994). Chang, T. C. (1987). “Network resource allocation using an expert system with

fuzzy logic reasoning.” Ph.D. thesis, University of California, Berkeley, CA. Dasgupta, G. (2008). “Stiffness matrix from isoparametric closed form shape

functions using exact integration.” J. Aerosp. Eng. in press. Duan, L., Loh, J. T., and Chen, W. F. (1990). “M-P-f-based analysis of dented

tubular members.” Struct. Engrg. Rep. No. CE-STR-90-27, School of Civ. Engrg., Purdue Univ., West Lafayette, Ind.

Duvant, G. and Lions, J. L. (1972). Les inéquations en méchanique et en physique.

Dunod, Paris, France (in French). Eshenaur, S. R., Kulicki, J. M., and Mertz, D. R. (1991). “Retrofitting

distortion-induced fatigue cracking of non-composite steel girder-floorbeam-stringer bridges.” Proc., 8th Annual Int. Bridge Conf., Engineers’ Soc. of Western Pennsylvania, Pittsburgh, Pa., 380–388.

Federal Register (1988). 33(No. 146; July 27), 10756. Federal Highway Administration (FHWA) (1991). Evaluating scour at bridges.

Rep., Hydr. Engrg. Circular No. 18: FHWA-IP-90-017, Washington, D.C. Frater, G. S. and Packer, J. A. (1992a). “Weldment design for RHS truss connec-

tions. I: Applications.” J. Struct. Engrg., ASCE, 118(10), 2784–2803. Frater, G. S. and Packer, J. A. (1992b). “Weldment design for RHS truss con-

nections. II: Experimentation.” J. Struct. Engrg., ASCE, 118(10), 2804–2820. Garrett, D. L. (2003). “Coupled analysis of floating production systems.” Proc.,

Int. Symp. on Deep Mooring Systems, (CD-ROM), ASCE, Reston, VA, 152– 167.

Gaspar, N. and Koenders, M. A. (2001a). “Estimates of the shear modulus of a granular assembly using heterogeneous media techniques.” Powders and Grains 2001, Y. Kishino, ed., A.A. Balkema, Lisse, 389–392.

Gaspar, N. and Koenders, M. A. (2001b). “Micromechanic formulation of macro-

scopic structures in a granular medium.” J. Engrg. Mech., 127(10), 987–993. Goossens, M., Mittlebach, F., and Samarin, A. (1994). The L^{A}TEX Companion.

Addison–Wesley Pub. Co., Reading, Mass. Huang, Y., Bird, R., and Bell, M. (2009a). “A comparative study of the emission

by road maintenance works and the disrupted traffic using life cycle assessment and micro-simulation.” Transportation Research Part D, 14, 197–204.

Huang, Y., Bird, R., and Hendrich, O. (2009b). “Development of a life cycle assessment tool for construction and maintenance of asphalt pavements.” Journal of Cleaner Production, 17, 283–296.

International Conference of Building Officials (1988). Uniform building code.

Whittier, Calif. Ireland, H. O. (1954). “Stability analysis of Congress Street open cut in Chicago.”

Géotechnique, London, England, 4(4), 163–168. Kuhn, M. R. (2011). “Style files for ASCE-like documents.” ascelike,

<http://www.ctan.org/tex-archive/macros/latex/contrib/ascelike> (Aug. 20, 2011).

Lamport, L. (1994). L^{A}TEX: A Document Prepartation System User’s Guide and

Reference Manual. Addison–Wesley Pub. Co., Reading, Mass. Liggett, J. A. and Caughey, D. A. (1998). “Fluid statics.” Fluid mechanics, (CD-

ROM), ASCE, Reston, VA, Chapter 2, 167–177. Lotus 1-2-3 reference manual; release 2.01 (1985). Lotus Development Corp.,

Cambridge, Mass. Melan, J. (1913). Theory of arches and suspension bridges. Myron C. Clark,

Chicago, Ill D. B. Steinman, translator. Moody’s municipal & government manual (1988). Moody’s Investors Service, New

York, N.Y. Pennoni, C. R. (1992). “Visioning: the future of civil engineering.” J. Profl. Issues

in Engrg. Education and Practice, ASCE, 118(3), 221–233. Reiffenstuhl, H. (1982). “Das vorspannen von bewehrung auf druck: Grundsaet-

zliches und anwendungsmoeglichkeiten [prestressing of reinforcing in compression: fundamentals and application possibilities].” Beton-und Stahlbetonbau, Berlin, Germany, 77(3), 69–73 (in German).

Sotiropulos, S. N. (1991). “Static response of bridge superstructures made of fiber

reinforced plastic.” M.S. thesis, West Virginia Univ., Morgantown, WV. Stahl, D. C., Wolfe, R. W., and Begel, M. (2004). “Improved analysis of timber

rivet connections.” J. Struct. Eng., 130(8), 1272–1279. Vesilind, P. A. (1992). “Discussion of ‘Guidance for engineering-design-class lec-

tures on ethics,’ by Richard H. McCuen.” J. Profl. Issues in Engrg. Education and Practice, ASCE, 118(2), 214–215.

Wichtmann, T. (2005). “Explicit accumulation model for non-cohesive soils under cyclic loading.” Ph.D. thesis, Institute for Soil Mechanics and Foundation Engineering, Ruhr-Univ. Bochum, Germany, , <www.rz.uni-karlsruhe.de/∼gn97/> (2007).

Zadeh, L. A. (1981). “Possibility theory and soft data analysis.” Mathematical frontiers of the social and policy sciences, L. Cobb and R. M. Thrall, eds., Westview Press, Inc., Boulder, Colo., 69–129.

APPENDIX I. NOTATION The following symbols are used in this paper:

D = pile diameter (m);
R = distance (m); and
C_{Oh no!} = fudge factor.